### v1

@@ 5 Convergence analysis
Now we analyze the convergence of Algorithm 1. We assume the Lipschitz continuity of the discounted payoff function $g_{k}(\cdot)$. {assumption} The discounted payoff function $g_{k}(\cdot)$ is $L$-Lipschitz continuous: for $k=0,1,\dots,N$
@@ 5.1 One-step error estimation
First, we analyze the one step error over the interval $[t_{k},t_{k+1}]$ for $k=0,1,\dots,N-1$, i.e., the numerical value function $\hat{u}_{k}(\mathbf{w})$ as an approximation to the exact one ${u}_{k}(\mathbf{w})$. Given the previous value function $\hat{u}_{k+1}$, we define the current CVF $\check{c}_{k}(\mathbf{W}_{k})$ by
@@
The classic backward Euler scheme for BSDE approximate $\nabla\check{c}_{k}$ by
@@
where $c_{k}^{\rm CLS}$ solves the continuous least squares problem Eq. 12 for $k=0,1,\dots,N-1$. In Theorem 5.4 below, we derive the convergence of $c_{k}^{\operatorname{CLS}}$ to the exact $c_{k}$ provided that the previous approximation error $\|u_{k+1}-\hat{u}_{k+1}\|_{L_{\omega_{k+1}}^{2}(\mathbb{R}^{d})}^{2}$, $\mathcal{E}_{k}^{\rm best}$ in Eq. 20, and time step size $\Delta t$ are small. We first provide in Lemma 5.1 an estimate of $\mathcal{E}_{k}^{\rm best}$ in Eq. 20, which is a direct application of [17, Theorem 4.2].
@@
For the hyperbolic cross index set $I$ with maximum order $p$ defined in Eq. 8, given $\check{c}_{k}\in\mathcal{K}^{m}(\mathbb{R}^{d},\omega_{k})$, we have
@@
Next, we provide an error estimate of solving the continuous least squares problem Eq. 12 in terms of the error of the previous value function, the best approximation error in the sparse Hermite polynomial ansatz space Eq. 20 and the time step size $\Delta t$. The proof is inspired by the foundational work [12].
@@
The equality Eq. 12 implies that $E_{k}(c_{k}^{\rm CLS})\leq E_{k}(\psi)$ for all $\psi\in P_{I,k}$. Taking $\psi:=c_{k}^{*}$ as defined in Eq. 19, we obtain
@@
where the last inequality follows from the condition $2\Delta t\leq 1$ and Lemma 5.2. This shows the estimate Eq. 23, and completes the proof.
@@ 5.2 Global error estimation
Theorem 5.8 implies that the numerical value function $\hat{u}_{k}$ computed by Algorithm 1 at each time $t_{k}$ approximates the exact one ${u}_{k}$ well if $\Delta t$, $\mathcal{E}_{k}^{\rm best}$, and $\mathcal{E}_{k}^{\rm stat}$ are small. Using Lemma 5.1, the best approximation error $\mathcal{E}_{k}^{\rm best}$ decays geometrically with the increase of the order $p$ of Hermite polynomials. By the law of large numbers, the statistical error $\mathcal{E}_{k}^{\rm stat}$ between the discrete and the continuous least squares will be small when using a large number of sample paths.
@@ 6 Numerical examples
In this section, we present several examples of high-dimensional Bermudan option pricing to show the efficiency and accuracy of Algorithm 1. The codes for the numerical experiments can be founded in the GitHub repository https://github.com/jiefeiy/glsm-american. The accuracy of the computed prices $\hat{v}_{0}(\mathbf{s}_{0})$ and deltas $\nabla\hat{v}_{0}(\mathbf{s}_{0})$ are measured by the relative errors defined by
@@
respectively, where $v_{0}^{\dagger}$ and $\Delta_{0}^{\dagger}$ are exact price and delta at time $t=0$. Unless otherwise stated, the results are the average of 10 independent runs. The parameter settings of the examples are summarized in Table 1, which has been considered previously [25, 21, 6, 3, 9].
@@ 6.2 Example 2: American geometric basket call
[…]
@@
Next, Fig. 2 shows the classification results of continued and exercised data using G-LSM and LSM with the number of simulated paths $M=100,000$ in $d=7$ or 20. Compared with the exact exercise boundary, G-LSM achieves better accuracy in determining the exercise boundary than LSM despite of using the same ansatz and number of paths. Thus, even with the right ansatz space, LSM might fail the task of finding exercise boundary in high dimensions using only a limited number of samples. Compared with [6, Figure 5] and [18, Figure 6], Fig. 2 domonstrates that G-LSM can detect accurate exercise boundary with fewer number of paths than the DNN-based method.
@@ 6.3 Example 3: Bermudan max-call with symmetric assets
Table 6 reports the time $T_{\rm lin}$ of solving linear system using the built-in MATLAB function cgs. The main memory requirement is storing the matrix $A\in\mathbb{R}^{M\times N_{b}}$, which takes $M\times N_{b}\times 8$ bytes in double-precision floating-point format. For $d=100$, the memory for storing $A$ is about 4GB. This limits the application of G-LSM with direct solvers in higher dimensions. To remedy the issue, one can solve the linear system on a large RAM server, or use single-precision floating-point, or with stochastic gradient descent.
Figure 4 shows the classification of continued and exercised sample points computed by G-LSM and LSM in the example of two-dimensional max-call. G-LSM yields a smoother exercise boundary than LSM. Compared with the exercise boundary computed in literature [19, Figure 3], G-LSM exhibits higher accuracy, albeit that the same ansatz space for the CVF is employed.
@@ 7 Conclusions and outlook
There are several avenues for further research. The superiority of G-LSM over LSM in high dimensions indicates that matching option values at $t_{k+1}$ might be a better choice than at $t_{k}$ for approximating the continuation value function. There are other variants of LSM with different ansatz spaces, and it is natural to ask whether incorporating the gradient information will also result in improved performance for these variants. Moreover, to solve higher dimensional problems, e.g., $d=1000$, the hierarchical tensor train technique can be applied, which has been combined with LSM in [2], and it is natural to combine the technique with G-LSM. Numerical results also show the potential of G-LSM for stochastic volatility models. It would be interesting to investigate G-LSM for more challenging financial models, e.g., rough volatility models.
@@ Acknowledgments
The authors acknowledge the support of research computing facilities offered by Information Technology Services, the University of Hong Kong.

### v2

@@ 5 Convergence analysis
Now we analyze the convergence of Algorithm 1. We restrict the theoretical analysis to the multi-asset Black-Scholes model presented in Section 2.2 since the corresponding CVF can be reformulated as a function of independent Brownian motions and the associated best approximation error in the sparse Hermite polynomial ansatz space $P_{I,k}$ can be rigorously established, cf. Lemma 5.1. In practice, Algorithm 1 can be applied to general models driven by Itô diffusions. See Section 6.5 for one numerical example related to the Heston model. We assume the Lipschitz continuity of the discounted payoff function $g_{k}(\cdot)$. {assumption} The discounted payoff function $g_{k}(\cdot)$ is $L$-Lipschitz continuous: for $k=0,1,\dots,N$
@@ 5.1 One-step error estimation
First, we analyze the one step error over the interval $t_{k},t_{k+1}]$ for $k=0,1,\dots,N-1$, i.e., the numerical value function $\hat{u}_{k}(\mathbf{w})$ as an approximation to the exact one ${u}_{k}(\mathbf{w})$. Given the previous value function $\hat{u}_{k+1}$, we define the current CVF $\check{c}_{k}(\mathbf{W}_{k})$ by
@@
The classic backward Euler scheme for BSDE approximates $\nabla\check{c}_{k}$ by
@@
where $c_{k}^{\rm CLS}$ solves the continuous least squares problem Eq. 12 for $k=0,1,\dots,N-1$. In Theorem 5.4 below, we derive the convergence of $c_{k}^{\operatorname{CLS}}$ to the exact $c_{k}$ provided that the previous approximation error $\|u_{k+1}-\hat{u}_{k+1}\|_{L_{\omega_{k+1}}^{2}(\mathbb{R}^{d})}^{2}$, $\mathcal{E}_{k}^{\rm best}$ in Eq. 28, and time step size $\Delta t$ are small. We first provide in Lemma 5.1 an estimate of $\mathcal{E}_{k}^{\rm best}$ in Eq. 28, which is a direct application of [20, Theorem 4.2].
@@
For the hyperbolic cross index set $I$ with maximum order $p$ defined in Eq. 8, given $\check{c}_{k}\in\mathcal{K}^{m}_{\omega_{k}}(\mathbb{R}^{d})$, we have
@@
Next, we provide an error estimate of solving the continuous least squares problem Eq. 12 in terms of the error of the previous value function, the best approximation error in the sparse Hermite polynomial ansatz space Eq. 28 and the time step size $\Delta t$. The proof is inspired by the foundational work [15]. We first derive a one-step error propagation by using Young’s inequality and then apply discrete Gronwall’s inequality to establish global convergence. In contrast to the approach in [15], our analysis also incorporates the statistical error introduced by the Monte Carlo method and the one-step error proven in Lemma 5.2. Moreover, the best approximation error of sparse Hermite polynomial approximation is given in Lemma 5.1.
@@
The equality Eq. 12 implies that $E_{k}(c_{k}^{\rm CLS})\leq E_{k}(\psi)$ for all $\psi\in P_{I,k}$. By taking $\psi:=c_{k}^{*}$ as defined in Eq. 27, we obtain
@@
where the last inequality follows from the condition $2\Delta t\leq 1$ and Lemma 5.2. This shows the estimate Eq. 31, and completes the proof of the theorem.
@@ 5.2 Global error estimation
Theorem 5.8 implies that the numerical value function $\hat{u}_{k}$ computed by Algorithm 1 at each time $t_{k}$ approximates the exact one ${u}_{k}$ well if $\Delta t$, $\mathcal{E}_{k}^{\rm best}$, and $\mathcal{E}_{k}^{\rm stat}$ are small. Using Lemma 5.1, the best approximation error $\mathcal{E}_{k}^{\rm best}$ decays geometrically with the increase of the order $p$ of sparse Hermite polynomials. By the law of large numbers, the statistical error $\mathcal{E}_{k}^{\rm stat}$ between the discrete and the continuous least squares will be small when using a large number of sample paths.
@@ 6 Numerical examples
In this section, we present several examples of high-dimensional Bermudan option pricing to show the efficiency and accuracy of Algorithm 1. In practice, the proposed method is applicable to other problems, e.g., high-dimensional optimal stopping problems, provided that the underlying dynamics follow an Itô process.
The codes for the numerical experiments can be founded in the GitHub repository https://github.com/jiefeiy/glsm-american. The accuracy of the computed prices $\hat{v}_{0}(\mathbf{s}_{0})$ and deltas $\nabla\hat{v}_{0}(\mathbf{s}_{0})$ are measured by the relative errors defined by
@@
respectively, where $v_{0}^{\dagger}$ and $\Delta_{0}^{\dagger}$ are exact price and delta at time $t=0$. Unless otherwise stated, the results are the average of 10 independent runs. The parameter settings of the examples are summarized in Table 1, which has been considered previously [31, 27, 6, 3, 9]. The least squares problem is solved using the built-in MATLAB function cgs. The maximum order $p$ of the hyperbolic cross index set is one important hyper-parameter. In the experiments, we gradually increase the maximum order such that the pricing results do not change significantly, following the standard practice in sparse polynomial approximation [21]. See Table 3 for the study on the pricing results with respect to the maximum order $p$. In practice, a higher-dimensional problem often needs a lower maximum order of hyperbolic cross index set, which might be related to the intrinsic low dimensional structure of high dimensional functions.
@@ 6.2 Example 2: American geometric basket call
Table 5: Deltas of American geometric basket call at $t=0$ using $M=720,000$ samples. $\mathbf{1}=(1,1,\dots,1)^{\top}$.
@@
Next, Fig. 2 shows the classification results of continued and exercised data using G-LSM and LSM with the number of simulated paths $M=100,000$ in $d=7$ or 20. Compared with the exact exercise boundary, G-LSM achieves better accuracy in determining the exercise boundary than LSM despite of using the same ansatz and number of paths. Thus, even with the right ansatz space, LSM might fail the task of finding exercise boundary in high dimensions using only a limited number of samples. Compared with [6, Figure 5] and [22, Figure 6], Fig. 2 domonstrates that G-LSM can detect an accurate exercise boundary with fewer number of paths than the DNN-based method.
@@ 6.3 Example 3: Bermudan max-call with symmetric assets
A method known as deep empirical risk minimization (deep ERM) has been introduced in [25]. This approach involves parameterizing the stopping boundary using a neural network and includes Example 3 in [25, Table 4]. A comparison between Table 6 and [25, Table 4] reveals that both methods are effective in handling high dimensions. The runtime of the G-LSM method is shorter than that of deep ERM for moderately high dimensions (e.g., $d=2,3,5,10$), but longer for higher dimensions (e.g., $d=20,50$). This difference can be attributed to the fact that deep neural network-based approaches are generally less sensitive to dimensions but require a larger number of paths for training. For instance, the results presented in [25, Table 4] were obtained using $1,536,000$ simulation paths, whereas Table 6 utilized only $100,000$ paths.
@@
We mention two bottlenecks of the proposed G-LSM for solving higher-dimensional problems. The first is the main memory requirement of the method, i.e., storing the matrix $A\in\mathbb{R}^{M\times N_{b}}$, which takes $M\times N_{b}\times 8$ bytes in double-precision floating-point format. For $d=100$, the memory for storing $A$ is about 4GB. This limits the application of G-LSM with direct solvers in higher dimensions. To remedy the issue, one can solve the linear system on a large RAM server, or use single-precision floating-point, or with stochastic gradient descent. Second, although hyperbolic cross type sparse index can overcome the curse of dimensionality in the sense that the computational complexity does not grow exponentially with respect to $d$, cf. Fig. 3, the practical implementation for large $d$ can still be non-trivial even when the computational complexity grows only polynomially with respect to $d$. Thus, we have presented the numerical examples only up to $d=100$.
Figure 4 shows the classification of continued and exercised sample points computed by G-LSM and LSM in the example of two-dimensional max-call. G-LSM yields a smoother exercise boundary than LSM. Compared with the exercise boundary computed in literature [24, Figure 3], G-LSM exhibits higher accuracy, albeit that the same ansatz space for the CVF is employed.
@@ 7 Conclusions and outlook
There are several avenues for further research. The superiority of G-LSM over LSM in high dimensions indicates that matching option values at $t_{k+1}$ might be a better choice than at $t_{k}$ for approximating the continuation value function. There are other variants of LSM with different ansatz spaces, and it is natural to ask whether incorporating the gradient information will also result in improved performance for these variants. Moreover, to solve higher dimensional problems, e.g., $d=1000$, the hierarchical tensor train technique can be applied, which has been combined with LSM in [2], and it is natural to combine the technique with G-LSM. Numerical results also show the potential of G-LSM for stochastic volatility models. It would be interesting to investigate both numerically and theoretically the proposed G-LSM for even more challenging financial models, e.g., rough volatility model. Numerically, for such non-Markovian models, the algorithm would require suitable Markovian approximation, and theoretically, one key aspect is to establish a suitable framework for analyzing the approximation error. Furthermore, it is of interest to assess the proposed G-LSM on practical applications, by a suitable combination with techniques for calibrating model parameters from real market data.
@@ Acknowledgments
The authors acknowledge the support of research computing facilities offered by Information Technology Services, the University of Hong Kong. The authors are also grateful to two anonymous referees for their constructive comments which have led a significant improvement in the quality of the paper.